OpenStudy (anonymous):

Add these two complex numbers : (2 + 3i) + (-3 - i)

OpenStudy (anonymous):

complex numbers act as variables so what this is \[(2 + 3i) + (-3-i)=(2-3)(3i-i)= -1+2i\]

OpenStudy (whpalmer4):

You wrote that middle part incorrectly: \[(2+3i) + (-3-i) = (2-3) + (3-1)i = -1 + 2i\]

OpenStudy (anonymous):

true thanks for catching that i've been having a bad day :(

OpenStudy (whpalmer4):

that means it can only get better, right? :-)

OpenStudy (anonymous):

until i hit \[- \infty\]i can only go up

OpenStudy (anonymous):

go down* see its getting worst lol

OpenStudy (anonymous):

lol :0 imma stop talking lol

OpenStudy (anonymous):

lol i saw that no reason to delete is XD

OpenStudy (whpalmer4):

anyhow, I just didn't want the OP to get confused and think there was a multiplication going on in the middle step there...

OpenStudy (anonymous):

yea no problem thanks for the fix